OpenStudy (anonymous):

When two atoms of 2H(deuterium) are fused to form one atom of 4He (helium), the total energy evolved is 3.38 x 10-12 joules. What is the total change in mass (in kilograms) for this reaction?

OpenStudy (anonymous):

you use einsteins equation: E=m*c^2 so mass is: m=E/c^2